OpenStudy (anonymous):

Oliver's grandfather invested $150 at 5% interest compounded quarterly. When Oliver inherited it, the amount was $6000. For how many years had the $150 been invested

OpenStudy (anonymous):

P = N(1+i)^n I believe is the formula.

OpenStudy (anonymous):

no it's A=P[1+(r/n)]^ny

OpenStudy (anonymous):

Yea same thing. Sub your givens into the formula. Isolate and solve for t.

OpenStudy (campbell_st):

is the interest 5% per quarter or 5% per annum...?

OpenStudy (anonymous):

quarter

OpenStudy (campbell_st):

ok then its 6000 = 150(1+0.05)^n divide both sides by 150 40 = 1.05^n take the log of both sides log(40) = nlog(1.05) n = log(40)/log(1.05) n = 75.607 divide by 4 you get the number of years